Chapter 20 DQ has two questions, Q1 and Q2, with multiple parts (a and b). Please answer completely. For Q2 you must show your work to receive points! If you just give the answer for Q2 you will not receive any points.

Q1. Briefly explain the (a) advantages and (b) disadvantages of unemployment insurance benefits.

Q2. Calculate the (a) Unemployment Rate and (b) the Labor Force Participation rate for the imaginary country of VanFeeney. VanFeeney has the labor force classification as follows:  Employed 80 million, Unemployed 20 million, Not in the Labor Force 50 million.
Answer #1

Advantages

  • Unemployment insurance benefits help individuals who have lost their job to sustain a desirable consumption level.
  • It ensures a fixed income till the person gets employment in order to fulfill basic needs.
  • The main benefits are insurance and incentives.

Disadvantages

  • Unemployed Individuals may become reluctant to find a job.
  • The benefits decrease over the period of unemployment spell.
  • The benefits are provided for restricted time period.
  • A tax is imposed after reemployment which increases with the length of the previous unemployment spell.

Unemployment Rate =( No. of People Unemployed / Total Labor Force ) * 100

Labor Force = Total Employed + Total Unemployed = 80 + 20 = 100

Unemployment Rate = ( 20 / 100 ) * 100 = 20%

Labor Force Participation Rate = (Labor Force / Total Adult Population) * 100

= (100 / 80 + 20 ) * 100

= (100 / 100) * 100 = 100%

In the absence of any information Total Population is assumed to be

Total Employed + Total Unemployed = 100 Million